OpenStudy (anonymous):

simplify the cube root of 25v^2w^2 (multiplied by) cube root 25vw

OpenStudy (anonymous):

\[\sqrt[3]{25v^2w^2}.\sqrt[3]{25vw}=\sqrt[3]{5^4v^3w^3}=(5vw)\sqrt[3]{5}\]